# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v2

@@ 3 Conclusion
In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that provides ready-to-use embeddings with worldwide coverage at 10-meter resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data. TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing. We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models. Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
@@ Acknowledgments
We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
@@
Paper review and refinement: AB, CA, DAC, SJ, SK
@@
The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) datasest [4].
@@
The AGB downstream task uses the open BioMassters dataset [111].
The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners, hence data access will require a Data Transfer Agreement.
@@ A.3 Time Series Extraction and d-pixel Definition
In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time. Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there is valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
@@ A.4 The d-pixel Representation
We represent each 10-m pixel in the time series of images from annual multi-spectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig. 1 a). In this array, each row corresponds to a chronologically ordered observation date, and each column is a specific spectral band or polarization. This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities. Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored during attention computation or ignored in the temporal sampling process.
@@ A.5 Dual-Encoder Architecture
  • Sentinel-2 MSI Encoder: This branch processes a time series of 10 spectral bands (excluding the 60 m bands used for the detection of water vapor and clouds) from Sentinel-2. We used blue (B2), green (B3), red (B4), red edges 1–3 (B5, B6, B7), near-infrared (B8, B8A), and shortwave infrared (B11, B12).
@@ A.6 Projector Network
The fused embedding from the dual-encoder stage is subsequently fed into a large projector network. This projector is a five-layer MLP, with each hidden layer having 16,384 dimensions. This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [187], to enable effective redundancy reduction during the self-supervised loss computation. The final output of the projector for each input d-pixel is an embedding, which is then used in the loss calculation. For downstream tasks, we use the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding. The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
@@ A.7.1 Augmented View Generation
    For each view, independently sampling a fixed number of valid observation dates from the annual Sentinel-2 time series (10 spectral bands).
@@
    For each view, independently sampling a fixed number of valid observation dates from the annual Sentinel-1 time series (2 polarizations).
@@ A.7.2 Loss Function
To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al. [16]. This involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating mixed views $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$, where $\alpha_{mix}\sim\text{Beta}(\beta_{p},\beta_{p})$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained. The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
@@
where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model’s outputs. The total loss function optimized during the training of TESSERA is a weighted sum:
@@ A.7.3 Pre-training Details
The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles. d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400, i.e, selecting one pixel from each 20$\times$20 patch. Each d-pixel contained the annual time series for 10 spectral bands of Sentinel-2 data and the annual time series of VV and VH polarizations for Sentinel-1 data.
@@
Learning Rate Schedule - Linear warmup (10%), plateau (20%), cosine decay
@@
Figure 7: TESSERA is trained on globally-distributed training data with greater data density in the northern hemisphere. TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024. The color of each hexagon in the map corresponds to the number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate color legend. This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule included a linear warming during the initial 10% of steps, a plateau for the next 20%, followed by a cosine decay. The global batch size was 32768.
@@
The impact of this approach on training stability is empirically demonstrated in Fig. 8b. Compared to a conventional, localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
Figure 8: Data shuffling greatly improves training stability. a, Schematic of the data shuffling and loading process. D-pixels (colored squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary [152] performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
@@ A.8 Global Embedding Map Generation
A primary output of the TESSERA project is the generation of annual global embedding maps with 10-meter resolution for the years 2017-2024. To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used. For each 10-meter pixel on the globe and for each year:
@@
    The full Sentinel-1 and Sentinel-2 time series data at 10-meter resolution are acquired and preprocessed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
@@
This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 10-meter grid.
@@ Appendix B Scaling with data and network size
In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning. We performed a parameter sweep over the projector’s depth (number of layers) and width (number of neurons per layer), with results shown in Fig. 9a. As discussed by Zbontar et al. [187], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task. The highest validation score was achieved with the largest configuration tested (8 layers deep, 32,768 neurons wide). However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure. For the final TESSERA v1.0 model, we selected the configuration with 4 layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
Furthermore, we analyzed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig. 9b. The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget. This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks. The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
Figure 9: Model performance improves with increasing projector size and additional computation. a, Hyperparameter sweep for the projector network’s architecture. Each point represents a configuration, with its position determined by depth (x-axis) and width (y-axis). The color and size of the point indicate the validation F1 score and the number of parameters, respectively. Performance generally increases with both width and depth. The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints. The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency. b, Scaling laws for TESSERA. The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters). A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used. The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
@@ Appendix C Ablation study
To validate the contribution of each key component within the TESSERA framework, we conducted an ablation study. We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
@@
The results of the ablation study are summarized in Table 4. Our analysis reveals several key insights into the model’s architecture and training strategy. The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches. Removing mixup regularization also severely degrades results. When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable, robust representations.
@@ Appendix D Downstream Task Application Methodology
A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labeled data. Having pre-trained TESSERA on large unlabeled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications. The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
@@
    Prepare Labeled Downstream Data: The labeled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
@@
    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA embeddings as input features and the corresponding labels. Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used. This training typically requires significantly less labeled data and computational power compared to training a deep model from scratch.

### v3

@@ 3 Conclusion
In this work, we introduce TESSERA, an innovative Remote Sensing Foundation Model that provides ready-to-use embeddings with worldwide coverage at 10 m resolution through self-supervised learning applied to pixel-level optical and microwave satellite time series data. TESSERA embeddings establish new performance standards, while our open-source methodology and GeoTessera library permit use of our work even by non-specialists who lack access to high-performance computing. We evaluated TESSERA’s capabilities across five distinct tasks, benchmarking our approach against leading task-specific models and existing foundation models. Our findings demonstrate that TESSERA surpasses both conventional RF baselines and current state-of-the-art geospatial foundation models in these diverse downstream applications.
@@ Acknowledgments
We gratefully acknowledge help from AMD Inc., Tarides, Jane Street, the Dawn supercomputing team at Cambridge, the Aalto University Science-IT project and the Aalto Scientfic Computing team, and the UKRI STFC AIRR Programme for providing us access to compute resources.
@@
Paper review and refinement: All
@@
The burn area detection task uses the open Monitoring Trends in Burn Severity (MTBS) dataset [4].
@@
The AGB downstream task uses the open BioMassters dataset [110]. The code for the winning model of the BioMassters competition is described in [110] and freely available on the corresponding GitHub page.
The implementation of SpectralGPT+UPerNet model is available in the Pangaea benchmark [100].
The stocking index downstream task uses plot data from 38 plots in Brazil. This contains private information regarding plot owners; hence, data access will require a Data Transfer Agreement.
@@ A.3 Time Series Extraction and d-pixel Definition
In other words, the d-pixel provides all spectral values (Sentinel-2) or backscatter values (Sentinel-1) at a given point over time. Note that d-pixels can be sparse and are accompanied by a mask vector $m_{i,j}$ of size $T$ that indicates the timesteps for which there are valid data, with a value 1 indicating that the corresponding row in $P_{i,j}$ is valid.
@@ A.4 The d-pixel Representation
We represent each 1- m pixel in the time series of images from annual multispectral (Sentinel-2) and SAR backscatter (Sentinel-1) as a “d-pixel” array (see Fig. 1 a). In this array, each row corresponds to a chronologically ordered observation date, and each column is a specific spectral band or polarization. This representation accommodates data gaps encountered in remote sensing due to cloud cover (for optical data) or other atmospheric interference and acquisition irregularities. Observations identified as invalid (e.g., via Sentinel-2 scene classification layers for clouds, or missing Sentinel-1 acquisitions) are masked and are handled by downstream model components, for instance, by being ignored in the temporal sampling process.
@@ A.5 Dual-Encoder Architecture
  • Sentinel-2 MSI Encoder: This branch processes a time series of 10 spectral bands (excluding the 60 m bands used for the detection of water vapour and clouds) from Sentinel-2. We used blue (B2), green (B3), red (B4), red edges 1–3 (B5, B6, B7), near-infrared (B8, B8A), and shortwave infrared (B11, B12).
@@ A.6 Projector Network
The fused embedding from the dual-encoder stage is subsequently fed into a large projector network. This projector is a six-layer MLP. Its architecture comprises an input layer mapping the fused embeddings to 16,384 dimensions, followed by four hidden layers of 16,384 dimensions each, and a final linear output layer. Each of the first five layers is a fully-connected linear layer followed by Batch Normalization and a ReLU non-linear activation function, making the network deep and highly non-linear. This significant expansion in dimensionality is crucial, as suggested by the original Barlow Twins work [189], to enable effective redundancy reduction during the self-supervised loss computation. The final output of the projector is then used in the loss calculation. For downstream tasks, we used the 128-dimensional output from the fusion MLP (before the projector) as the final pixel embedding. The TESSERA encoder (up to the fusion MLP) has approximately 46 million parameters, while the projector accounts for the majority of the model’s $\sim$1.4 billion parameters.
@@ A.7.1 Augmented View Generation
    For each view, independent sampling of a fixed number of valid observation dates from the annual Sentinel-2 time series (10 spectral bands).
@@
    For each view, independent sampling of a fixed number of valid observation dates from the annual Sentinel-1 time series (2 polarizations).
@@ A.7.2 Loss Function
To further enhance the robustness of the model and mitigate overfitting, TESSERA incorporates an additional mix-up regularization term $\mathcal{L}_{MIX}$, inspired by Bandara et al. [17]. For each training batch, this involves shuffling one set of views (e.g., $Y_{B}$) along the batch dimension to create $Y_{S}=\text{Shuffle}(Y_{B})$, then generating a mixed view $Y_{M}=\alpha_{mix}Y_{A}+(1-\alpha_{mix})Y_{S}$. The mixing coefficient $\alpha_{mix}$ is sampled from a uniform distribution on the unit interval, $\alpha_{mix}\sim U(0,1)$. The embeddings $Z_{M}$ and $Z_{S}$ are obtained from their respective views. The mix-up loss penalizes deviations from the assumption that a linear interpolation in the input space corresponds to a linear interpolation in the embedding space.
@@
where $C^{MA}=(Z_{M})^{T}Z_{A}$ and $C^{MS}=(Z_{M})^{T}Z_{S}$ are the actual cross-correlation matrices from the model’s outputs. The total loss function optimized during training is a weighted sum:
@@ A.7.3 Pre-training Details
The TESSERA model was pre-trained using approximately 800 million d-pixels drawn from 3,012 globally distributed Military Grid Reference System (MGRS) tiles. d-pixels were generated by spatially downsampling Sentinel-1 and Sentinel-2 data by a factor of 400. This was achieved by systematically taking every 20th pixel along the horizontal and vertical axes. Each d-pixel contained the annual time series for 10 spectral bands of Sentinel-2 data and the annual time series of VV and VH polarizations for Sentinel-1 data.
@@
Learning Rate Schedule - Linear warmup (10%) followed by cosine decay
@@
Figure 8: TESSERA is trained on globally-distributed training data. TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024. The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend. This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
As we were in a compute-limited rather than data-limited pre-training regime, we trained for a single epoch, which corresponded to approximately 6200 GPU hours on 16 AMD MI300X GPUs (192GB memory each). We used PyTorch with Fully Sharded Data Parallel (FSDP) and Automatic Mixed Precision (AMP) enabled. AdamW optimizer was used with a base learning rate of 0.002 and weight decay of $1\times 10^{-6}$. The learning rate schedule consisted of a linear warmup during the initial 10% of steps, followed immediately by a cosine decay for the remainder of training. The global batch size was 32768. Key training dynamics, including loss curves, learning rate schedule, and evolution of downstream performance during pre-training, are visualized in Supplementary Fig. 9.
Figure 9: Training dynamics curves show that TESSERA rapidly reaches its asymptotes. a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss. b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay. c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction. d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant. e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
@@
The impact of this approach on training stability is empirically demonstrated in Fig. 10b. Compared to a conventional localized shuffling strategy, which results in a highly volatile loss curve (top plot), our global shuffling strategy yields a markedly smoother and more stable convergence (bottom plot). This enhanced stability is fundamental for robust convergence and for preventing the model from overfitting to regional characteristics.
Figure 10: Data shuffling greatly improves training stability. a, Schematic of the data shuffling and loading process. D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary [154] performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
@@ A.8 Global Embedding Map Generation
A primary output of the TESSERA project is the generation of annual global embedding maps with 1- meter resolution for the years 2017-2024. To generate these maps, the pre-trained and frozen TESSERA dual encoder (excluding the projector) is used. For each 1- meter pixel on the globe and for each year:
@@
    The full Sentinel-1 and Sentinel-2 time series data at 1- meter resolution are acquired and pre-processed to form d-pixels. Unlike pre-training, no spatial downsampling is performed at this stage.
@@
This process is repeated for all land pixels globally to create an annual embeddings map of shape (H, W, 128), where H and W are the dimensions of the global 1- meter grid.
@@ Appendix B Scaling with data and network size
In contrast, the architecture of the projector network proved to be critical for effective self-supervised learning. As detailed in Section A.6, the full projector is a six-layer MLP, consisting of an input and an output layer for accommodating dimensional changes, and four core hidden layers. Our parameter sweep focused on this computationally intensive hidden block. We varied the number of these hidden layers (referred to as ’depth’ in Fig. 11a) and their width (number of neurons per layer), with results shown in Fig. 11a. As discussed by Zbontar et al. [189], deeper and wider projectors consistently improved performance, as measured by the validation F1 score in the downstream task. The highest validation score was achieved with the largest configuration tested (8 hidden layers, 32,768 neurons wide). However, this configuration led to an out-of-memory error with our standard global batch size; therefore, we halved the batch size to 4096 to enable stable training for this specific setup, indicated by the dagger symbol (†) in the figure. For the final TESSERA v1.0 model, we selected the configuration with four hidden layers and a width of 16,384, marked with a star (*), as it provided a strong balance between high performance and computational efficiency.
Furthermore, we analysed the relationship between model performance, the number of trainable parameters, and the total training computation (FLOPs), as shown in Fig. 11b. The results demonstrate a clear scaling trend: model performance improves predictably with increases in model size and computational budget. This confirms that TESSERA’s performance can be further enhanced with greater computational resources, following established scaling laws for large neural networks. The final TESSERA v1.0 model is marked with a star (*) on this curve, highlighting its position as a model that balances state-of-the-art performance with cost-effective training requirements.
Figure 11: Model performance improves with increasing projector size and additional computation. a, Hyperparameter sweep for the projector network’s architecture. Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis). The colour and size of the point indicate the validation F1 score and the number of parameters, respectively. Performance generally increases with both width and depth. The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints. The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency. b, Scaling laws for TESSERA. The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters). A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used. The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
@@ Appendix C Ablation study
To validate the contribution of each key component within the TESSERA framework, we performed an ablation study. We systematically removed or disabled individual components from our full model (the baseline) and evaluated the impact on the 100x downsampled Austrian crop classification dataset.
@@
The results of the ablation study are summarized in Table 4. Our analysis reveals several key insights into the model’s architecture and training strategy. The most critical component turns out to be the global data shuffling pipeline, which is designed to break spatial autocorrelation and present the model with geographically diverse mini-batches. Removing mixup regularization also severely degrades results. When both are removed, the model’s performance collapses, showing that the combination of global shuffling and mixup regularization is needed to prevent overfitting and to learn generalizable robust representations.
@@ Appendix D Downstream Task Application Methodology
A core motivation for self-supervised learning with foundation models is the creation of task-agnostic feature embeddings that can be effectively transferred to various downstream tasks, particularly in scenarios with limited labelled data. Having pre-trained TESSERA on large unlabelled datasets, we evaluated the utility of its learned embeddings across different remote sensing applications. The evaluation methodology involves using pre-trained TESSERA encoders as fixed feature extractors.
@@
    Prepare labelled Downstream Data: The labelled dataset for the target task (e.g., pixel-level crop-type labels, canopy height measurements, or land use change polygons) is prepared.
@@
    Train Downstream Head: Only the parameters of this newly defined task head are trained using the extracted TESSERA embeddings as input features and the corresponding labels. Standard supervised learning techniques, optimizers (e.g., Adam), and task-appropriate loss functions (e.g., Cross-Entropy for classification, Mean Squared Error for regression) are used. This training typically requires significantly less labelled data and computational power compared to training a deep model from scratch.